An insulated tank initially contains 0.25 kg of a gas with an internal energy of 200 KJ/kg. Additional gas with an internal energy of 300 KJ/kg and an enthalpy of 400 KJ/kg enters the tank until the total mass of the gas contained is 1 kg. The final internal energy (in KJ/kg) of the gas in the tank is ?
a) 250
b) 275
c) 350
d) None of these

Please provide explaination to answer
1
Expert's answer
2019-08-27T09:31:57-0400

It is not closed system and we can not use adiabatic mixing. Enthalpy of additional gas will be converted to internal energy


"U=u_1 m_1+(1-m_1)h""U=200\cdot 0.25+(1-0.25)\cdot 400=350 (kJ)"

specific internal energy is


"u=350 (kJ\/kg)"